Ruel can paint a room in 5 hours anf vergel can do it in 6 hours.what part of the job can they do by working for 2 hours
Answer by nerdybill(7384)

Ruel can paint a room in 5 hours anf vergel can do it in 6 hours.what part of the job can they do by working for 2 hours
.
Let x = part
then
2(1/5 + 1/6) = x
2(6/30 + 5/30) = x
2(11/30) = x
11/15 = x
.
Answer:
11/15